Electromagnetic radiation of 5.44×10-11 m wavelength is shined on a metal surface and caused electron emission. Determine the work function of the metal if the maximum kinetic energy (Ek) of the emitted electron is 6.55×10-16 J.
Answer #1

Given:

lambda = 5.44*10^-11 m

use:

E = h*c/lambda

=(6.626*10^-34 J.s)*(3.0*10^8 m/s)/(5.44*10^-11 m)

= 3.654*10^-15 J

Given:

KE = 6.55*10^-16 J

Use:

KE = E - Eo

6.55*10^-16 J = 3.654*10^-15 J - Eo

Eo = 3.00*10^-15 J

Answer: 3.00*10^-15 J